### v1

@@ 4.2 $M_{3}$ is not a priority-neutral lattice
Figure 2 displays a lattice, typically denoted by $M_{3}$, which is a famous simple example in the lattice theory literature [Bir40, Grä12]. We use the facts proven in Section 4.1 to prove that $M_{3}$ never arises as the lattice of priority-neutral matchings.
@@
Figure 2: $M_{3}$ lattice and key steps of the proof of Theorem 4.8.
Theorem 4.8 ($M_{3}$ is not a $\mathsf{PN}$ lattice).
@@
Suppose for contradiction there is some set of preferences by which $\mathsf{PN}$ is isomorphic to $M_{3}$.
Using Theorem 4.7, consider the partial order $R$ on rotations which represents $\mathsf{K}$ (as defined above), and let $[\alpha_{1},\alpha_{2},\ldots,\alpha_{N}]$ be as in Theorem 4.7. Observe that this means there is some element $\mu_{1}\in\mathsf{PN}$ which corresponds to $\{\alpha_{1}\}\subseteq R$.
@@
However, by the definition of $M_{3}$, we know that $\mu_{1}\vee_{\mathsf{PN}}\mu_{s}=\mu_{1}\vee_{\mathsf{PN}}\mu_{t}$. Thus, we must have $\{\alpha\}\cup S=\{\alpha\}\cup T$. (For an illustration of this argument, see Figure 2.) This implies that $S=T$141414 This step of the proof would fail if we were not able to establish that at least one matching in $\mathsf{PN}$ differed from $\mathsf{IPDA}$ by only one rotation. Indeed, if we replaced $\{\alpha_{1}\}$ with a set of just two rotations $\{\alpha_{1},\beta\}$, then we would not be able to conclude from $\{\alpha_{1},\beta\}\cup S=\{\alpha_{1},\beta\}\cup T$ that $S=T$. This shows how we use our structural result Theorem 4.7. But this contradicts the assumption that $S$ and $T$ correspond to distinct elements of the $\mathsf{PN}$ lattice. This concludes the proof. ∎

### v2

@@ 4.3.3 Example of a Non-$\mathsf{PN}$ Lattice Due to the Good Chain Property
Figure 7 displays a lattice, typically denoted by $M_{3}$, which is a famous simple example of a non-distributive lattice from the order theory literature [Bir40, Grä12].
@@
Figure 7: $M_{3}$ lattice and key steps of the proof of 4.17.
Proposition 4.17 ($M_{3}$ is not a $\mathsf{PN}$ lattice).
@@
Suppose for contradiction there is some set of preferences by which $\mathsf{PN}$ is isomorphic to $M_{3}$. Let $G=(\mathsf{IPDA},R,E)$ be the fanout multigraph representing $\mathsf{PN}$, as given by 4.15. Let $\alpha_{1},\alpha_{2},\ldots,\alpha_{N}]$ be the ordering on $R$ as given by the good chain property. Observe that this means there is some element $\mu_{1}\in\mathsf{PN}$ which corresponds to $\{\alpha_{1}\}\subseteq R$.
@@
However, by the definition of $M_{3}$, we know that $\mu_{1}\vee_{\mathsf{PN}}\mu_{s}=\mu_{1}\vee_{\mathsf{PN}}\mu_{t}$. Thus, we must have $\{\alpha\}\cup S=\{\alpha\}\cup T$. (For an illustration of this argument, see Figure 7.) This implies that S=TS=T. [^22] This step of the proof would fail if we were not able to establish that at least one matching in $\mathsf{PN}$ differed from $\mathsf{IPDA}$ by only one rotation. Indeed, if we replaced $\{\alpha_{1}\}$ with a set of just two rotations $\{\alpha_{1},\beta\}$, then we would not be able to conclude from $\{\alpha_{1},\beta\}\cup S=\{\alpha_{1},\beta\}\cup T$ that $S=T$. This shows how we use the good chain property. But this contradicts the assumption that $S$ and $T$ correspond to distinct elements of the $\mathsf{PN}$ lattice. This concludes the proof. ∎